the domain and range of f(x)= 2x+1/x-1


1
Expert's answer
2021-11-17T00:21:26-0500

Solution;

"f(x)=\frac{2x+1}{x-1}"

Domain;

Means "x-1\neq0" hence,"x\neq1"

The domain of f(x) is all real values expect -1.

Range;

Let "f(x)=\frac{2x+1}{x-1}=y"

"2x+1=y(x-1)"

"2x+1=yx-y"

"2x-yx=-1-y"

"x(2-y)=-1-y"

"x=\frac{-1-y}{2-y}"

Means that the quantity "2-y\neq0"

Hence;

"y\neq2"

The range of f(x) is all real values except 2.